OpenStudy (anonymous):

Express the surface x = 5cosθsinφ, y = 6sinθsinφ, z = 4cosφ as a level surface f(x,y,z) = 14400, f(x,y,z) =

OpenStudy (anonymous):

Do you think you could show us any progress you have made to this point?

OpenStudy (anonymous):

x^2/(5sinφ)^2 + y^2/(6sinφ)^2 =1 (sinφ)^2 =1 - (z/4)^2 substitute for sinφ

OpenStudy (anonymous):

x^2/25 + y^2/36 = 1 - (z/4)^2 x^2/25 + y^2/36 + z^2/16 = 1

OpenStudy (anonymous):

multiply by 14400 on both sides

OpenStudy (anonymous):

576x^2 + 400y^2 + 900z^2 =14400= f(x,y,z)